OpenStudy (anonymous):

help what is Of the following sets, which numbers in {0, 1, 2, 3, 4} make the inequality 7x + 3 < 17 true

OpenStudy (anonymous):

please help

OpenStudy (dani_rose):

To make the inequality true you would need to plug in your opptions

OpenStudy (dani_rose):

\[(7\times2)+3<17\]

OpenStudy (dani_rose):

\[14+3<17\]

OpenStudy (anonymous):

{0} {0, 1} {0, 1, 2} {2, 3, 4}

OpenStudy (dani_rose):

\[17<17\]

OpenStudy (dani_rose):

2 is your answer

OpenStudy (anonymous):

which one a b c or d

OpenStudy (dani_rose):

I would say c.

OpenStudy (anonymous):

{0} {0, 1} {0, 1, 2} {2, 3, 4}

OpenStudy (dani_rose):

{0,1,2} Make the inequality true

OpenStudy (anonymous):

can u help with a few more

OpenStudy (dani_rose):

Sure

OpenStudy (anonymous):

Which of the following expressions can be used to find the area of a square with a side length of fraction 1 over 6 m?

OpenStudy (dani_rose):

Is there a pic?

OpenStudy (anonymous):

fraction 1 over 6 multiplied by 2 m2 2 multiplied by fraction 1 over 6 m2 fraction 1 over 6 whole squared m2 fraction 1 over 6 whole cubed m2

OpenStudy (dani_rose):

\[\frac{ 1 }{ 6 }\times2m^2\]

OpenStudy (dani_rose):

Is that what it's supposed to look like?

OpenStudy (anonymous):

yeah

OpenStudy (dani_rose):

\[2\times \frac{ 1 }{ 6 }m^2\]

OpenStudy (anonymous):

What is the area, in square units, of the polygon shown here?

OpenStudy (dani_rose):

A=a^2

OpenStudy (dani_rose):

Thats the formula

OpenStudy (dani_rose):

It's not letting me see the pic

OpenStudy (dani_rose):

I really gtg hopefully someone can help you

OpenStudy (anonymous):

someone help me

OpenStudy (anonymous):

anyone here